Write the null hypothesis in words and in symbols.

1. A random sample of 200 students got a mean score of 62 with s-5 in a knowledge test in Mathematics. In the standardization of the test, -50 with o-10.


1
Expert's answer
2021-06-22T12:33:29-0400

a.) The null hypothesis is a typical statistical theory which suggests that no statistical relationship and significance exists in a set of given single observed variable, between two sets of observed data and measured phenomena.

b.)

"n= 200"

"s=5"

"\mu=50"

"\sigma=10"

Using the null hypothesis,

"H_0 : \mu = 50"

"H_1 : \mu \ne 50"

Value of "z" can be calculated as "z = \dfrac{\bar x - \mu }{10} = \dfrac{62-50}{10} = 1.2"

At 0.05 level of significance "z_c= 1.96".


Hence calculated "|z|=1.2<1.96=z_c." Therefore null hypothesis is not rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 50, at the "\alpha=0.05"  significance level.